HUBBLE SPACE TELESCOPE DAILY REPORT #2164 PERIOD COVERED: 0000Z (UTC) 07/13/98 - 0000Z (UTC) 07/14/98 Daily Status Report as of 195/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera Two) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.2 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star HD149661. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.4 Completed NIC/1/2/3 7868 (Near-IR Cores of Radio Galaxies, Are the AGN's Moving in the Galaxy?) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the core region of radio galaxy M-84. Near-IR colors maps will determine the location of dense gas responsible for the obscuration and fueling of the AGNs. The observations were completed as scheduled, and no problems were reported. 1.5 Completed Four Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.6 Completed Six Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD 7951 (Anomalous Scatter Calibration) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star HD11067 without the use of an occulting bar in a long slit to make an anomalous scatter calibration observations. A wavelength calibration was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Twelve Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.9 Completed FOC/96 6618 (A Hidden Power Source in the QSO PG 1630+377) The Faint Object Camera (f/96) was used to make observations of quasar PG1630+337 in order to study its hidden power source. For this hidden source to be so bright as seen only through scattering indicates that it may be intrinsically the most significant power source shortward of the Lyman edge. An internal flat calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted 1.10 Completed STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.11 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed STIS/CCD/MA1 7288 (High Resolution Ultraviolet Spectroscopy of Hercules X-1/HZ Herculis) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make ultraviolet spectrographic observations of the X-ray binary pulsar 4U1656+35 (a.k.a., Hercules X-1/HZ Herculis). A wave length calibration observations was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.14 Completed NIC/1 7225 (Proper Motion of Stars Near SGR-A*) The Near Infrared Camera (Camera One) was to be used to make astrometric proper motion observations of stars in the vicinity of Sagittarius-A (a.k.a., NGC-4147). Sgr-A may be a black hole with a mass close to a million solar masses. As described in 2.1 and HSTAR 6584, the guide star acquisition for this proposal failed when the search radius limit was exceeded on FGS-2. This immediately followed the full maneuver update failures described in 2.1 and in HSTAR 6583. Four observations were affected when the take data flag remained down. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 6 HSTAR 6584 describes a guide star acquisition failure at 194/232222Z that occurred when the search radius limit was exceeded on FGS-2, immeediately following an FHST full maneuver update failure (HSTAR 6583). The proposal described in 1.14 was affected, as were an iteration of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) and STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) which appear during the interval of tomorrow's report. Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 10 Successful : 9 HSTAR 6583 documents FHST full maneuver updates which failed at 194/225957Z and at 194/230117Z following type T2 slews. These events immediately preceeded the acquisition failure described in 2.1. 2.3 Operations Notes: The SSR EDAC error counter was cleared three times using ROP SR-1. Following the acquisition failure (see 2.1), an attitude reference update was uplinked at 195/0044Z. At 195/0339Z, all six ephermeris table were uplinked. A DF-224 SPC dump to CCS string A was successfully performed at 195/0516Z 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH